PULSED-FIELD IONIZATION ELECTRON SPECTROSCOPY OF COPPER-PYRAZINE, -PYRIMIDINE, AND -PYRIDAZINE COMPLEXES

\underline{XU WANG}, JUNG SUP LEE, DONG-SHENG YANG, Department of Chemistry, University of Kentucky, Lexington, KY 40605-0055.

The copper complexes of pyrazine (1,4-C_{4}H_{4}N_{2}), pyrimidine (1,3-C_{4}H_{4}N_{2}), and pyridazine (1,2-C_{4}H_{4}N_{2}) are produced in laser-vaporization supersonic molecular beams and studied with pulsed field ionization zero electron kinetic energy spectroscopy and ab initio calculations. Adiabatic ionization energies of Cu-pyrazine, -pyrimidine and -pyridazine are measured to be 46038(7), 45332(7) and 43054(6) cm^{−1}, respectively. The metal-ligand stretching frequencies of Cu^{+}/Cupyrazine, -pyrimidine, and -pyrazine are 265/190, 270/189, and 273/192 cm^{−1}, and the metal-ligand bending frequencies of these complexes are 115/74, 161/151, and 131/123 cm^{−1}, respectively. In addition, a symmetric ring stretch of 906 cm^{−1} is observed for the Cu^{+}-pyrazine ion. In all three complexes, Cu binds with one nitrogen atom and is located in the same plane as the ligand ring.